# Three Scores and 15 Years (1948-2023) of Rao’s Score Test: A Brief History

### v1

@@ Abstract
This is the same modification suggested by Koenker, (1981). Note that the modification amounts to replacing $Var(\epsilon^{2}_{i})=\mu_{4}-\mu^{2}_{2}=3\sigma^{4}-\sigma^{4}=2\sigma^{4}$ (derived under normality) by a robust estimate, namely by, $\frac{1}{n}\sum^{n}_{i=1}(\widetilde{\epsilon}_{i}^{2}-\widetilde{\sigma}^{2})^{2}=({\nu^{\prime}\nu})/{n}$. For other applications of $RS^{*}(D)$ see for instance, Lucas, (1998) and Premaratne and Bera, (2017).
@@
Here the notation $RS^{*}(P)$ is used to signify robust RS test statistic under parametetric misspecification. Under $H_{0}:\psi=0$, $RS^{*}_{\psi}(P)$ asymptotically distributed as central $\chi^{2}_{r}$, i.e., $RS^{*}_{\psi}(P)$ has the same asymptotic distribution as of $RS_{\psi}$ in (69) based on the correct specification, thereby producing an asymptotically correct-size test under the locally misspecified alternative $l_{2}(\gamma,\phi)$.
$RS^{*}_{\psi}(P)$ essentially adjusts the asymptotic mean and variance of standard (unadjusted) $RS_{\psi}$. Another way to look at $RS^{*}_{\psi}(P)$ is to view the quantity, $\mathcal{I}_{\psi\phi\cdot\gamma}(\widetilde{\theta})\mathcal{I}_{\phi\cdot\gamma}(\widetilde{\theta})^{-1}S_{\phi}(\widetilde{\theta})$ as the prediction of $S_{\psi}(\widetilde{\theta})$ by $S_{\phi}(\widetilde{\theta})$. Here $S_{\phi}(\widetilde{\theta})$ is the score function of the parameter vector $\phi$ whose effect we want to take into account in constructing the robust version of the test. Therefore, $S_{\psi}(\widetilde{\theta})-\mathcal{I}_{\psi\phi\cdot\gamma}(\widetilde{\theta})\mathcal{I}^{-1}_{\phi\cdot\gamma}(\widetilde{\theta})S_{\phi}(\widetilde{\theta})$ is the part of $S_{\psi}(\widetilde{\theta})$ that remains after eliminating the effect of $S_{\phi}(\widetilde{\theta})$. Three more things regarding $RS^{*}_{\psi}(P)$ are worth noting. First, $RS^{*}_{\psi}(P)$ requires estimation only under the joint null, namely for the constrained model in which both $\psi=0$ and $\phi=0$. Given the full specification of the model $l(\gamma,\psi,\phi)$, it is of course possible to derive a RS test for $H_{0}:\psi=0$ in the presence of $\phi$. However, that requires MLE of $\phi$ which could be difficult to obtain in some cases. Second, when $\mathcal{I}_{\psi\phi\cdot\gamma}=0$, $RS^{*}_{\psi}(P)=RS_{\psi}$. This is a simple condition to check in practice. As mentioned earlier, if this condition is true, $RS_{\psi}$ is an asymptotically valid test in the local presence of $\phi$. Finally, Bera and Yoon, (1993) showed that for local misspecification $RS^{*}_{\psi}(P)$ is asymptotically equivalent to Neyman, (1959) $C(\alpha)$ test, and therefore, shares its optimality properties.
Example 5.4 To illustrate the usefulness of the robust score statistic $RS^{*}_{\psi}(P)$, we now consider the tests developed in Anselin et al., (1996) for the mixed regressive - spatial antiregressive (SAR) model with a SAR disturbance
@@
where $\widetilde{u}=y-X\widetilde{\gamma}$ are the OLS residuals, $\widetilde{\sigma}^{2}={\widetilde{u}^{\prime}\widetilde{u}}/{n}$ and $T=tr[(W^{\prime}+W)W]$. One very interesting observation here is that $RS_{\psi}$ is essentially same as the widely used Moran, (1948) I-test. Let us now consider testing $H_{0}$ under the local presence of $\phi$. First, the crucial quantity consider is $\mathcal{I}_{\psi\phi\cdot\gamma}$ which is equal to T and that can never be zero. Therefore robustification of $RS_{\psi}$ is needed. Anselin et al., (1996) derived the robust test as
@@
In a similar way we can find $RS_{\phi}$ and $RS^{*}_{\phi}(P)$ which are given by, respectively,
@@
This statistic be asymptotically distributed $\chi^{2}_{2}$. It is easy to verify that [see Bera et al., (2020, Corollary 1)]
@@
Anselin and Florax, (1995) and Anselin et al., (1996) provided simulation results on the finite sample performance of the unadjusted and adjusted RS tests and the related tests. The adjusted tests $RS^{*}_{\psi}(P)$ and $RS^{*}_{\phi}(P)$ performed remarkably well. Those had very reasonable empirical sizes, remaining within the confidence intervals in all cases. In terms of power they performed exactly the way they were supposed to.
@@
Now we combine the results of Sections 5.1 and 5.2 and develop robust tests $RS^{*}_{\psi}(DP)$ which will provide a two-way protection against both types of misspecifications, distributional $(D)$ and parametric $(P)$. As we have noted in (59), $RS^{*}_{\psi}(D)$ will involve both the $J(\theta)$ and $K(\theta)$ matrices in the variance expression of $S_{\psi}(\tilde{\theta})$. While to account of the parametric misspecification, as we did in (71), $\mathcal{I}_{\psi\phi\cdot\gamma}(\widetilde{\theta})\mathcal{I}^{-1}_{\phi\cdot\gamma}(\widetilde{\theta})S_{\phi}(\widetilde{\theta})$ must be subtracted from $S_{\psi}(\tilde{\theta})$ to center its mean to zero. The expression for $RS^{*}_{\psi}(DP)$ is given by [for details see Bera et al., (2020)]:
@@
We started this survey paper by stating that C.R. Rao’s work was always inspired by some practical problems. In his 2003 Econometric Theory (ET) Interview [see Bera, (2003, p.349)], on the RS test, Rao had the following to say, “The test evolved in a natural way while I was analyzing some genetic data. As I recall, the problem was the estimation of a linkage parameter using data sets from different experiments designed in such a way that each data set had information on the same linkage parameter. It was, however, necessary to test whether such an assumption could be made because of unforeseen factors affecting the experiments. This required a test for consistency of estimates derived from different experimental data sets.” Thus we had a new statistical test principle, after LR and W, motivated by a practical problem in genetics. However, as we have narrated here, the resulting RS test principle has a far reaching influence even beyond the Statistics, and in particular, it has become entrenced in the Econometrics literature as one of the most useful model misspecification testing tools. Apart from all the applications, of basing statistical inference, both in estimation and testing, the score function will remain useful in forseeable future, as discussed in Bera and Bilias, (2001). In the history of any scientific field, once in a while there comes a moment for a major breakthrough. Appearance of Rao, (1948) was such a historical moment. In fact, after that we have not witnessed any new test principle, beyond the trinity, LR, W and RS.
@@ Acknowledgements
We are indeed grateful to our research assistant (RA) Anirudh Adhikary. Without his assistance this paper wouldn’t have taken off. We are also grateful to our other RAs, Rong Yuwen, Tiancheng Guo, Scarlett He, Yice Zhang and Wenqi Zeng for their painstaking work on compiling the citation numbers and a very careful analysis. We wish we could report all the results they obtained. Thanks are also due to Dr. Malabika Koley for comments on an earlier draft, that improved the exposition of the paper. We however, retain the responsibility of any remaining errors. Part of this work was completed when the second author visited University of Illinois at Urbana-Champaign (UIUC).

### v2

@@ Abstract
This is the same modification suggested by Koenker, (1981). Note that the modification amounts to replacing $Var(\epsilon^{2}_{i})=\mu_{4}-\mu^{2}_{2}=3\sigma^{4}-\sigma^{4}=2\sigma^{4}$ (derived under normality) by a robust estimate, namely, by, $\frac{1}{n}\sum^{n}_{i=1}(\widetilde{\epsilon}_{i}^{2}-\widetilde{\sigma}^{2})^{2}=({\nu^{\prime}\nu})/{n}$. For other applications of $RS^{*}(D)$ see for instance, Lucas, (1998) and Premaratne and Bera, (2017).
@@
Here the notation $RS^{*}(P)$ is used to signify robust RS test statistic under parametetric misspecification. Under $H_{0}:\psi=0$, $RS^{*}_{\psi}(P)$ is asymptotically distributed as central $\chi^{2}_{r}$, i.e., $RS^{*}_{\psi}(P)$ has the same asymptotic distribution as of $RS_{\psi}$ in (69) based on the correct specification. Thus, $RS^{*}_{\psi}(P)$ provides an asymptotically correct-size test under the locally misspecified alternative $l_{2}(\gamma,\phi)$.
$RS^{*}_{\psi}(P)$ essentially adjusts the asymptotic mean and variance of standard (unadjusted) $RS_{\psi}$. Another way to look at $RS^{*}_{\psi}(P)$ is to view the quantity, $\mathcal{I}_{\psi\phi\cdot\gamma}(\widetilde{\theta})\mathcal{I}_{\phi\cdot\gamma}(\widetilde{\theta})^{-1}S_{\phi}(\widetilde{\theta})$ as the prediction of $S_{\psi}(\widetilde{\theta})$ by $S_{\phi}(\widetilde{\theta})$. Here $S_{\phi}(\widetilde{\theta})$ is the score function of the parameter vector $\phi$ whose effect we want to take into account in constructing the robust version of the test. Therefore, the net score $S^{*}_{\psi}(\widetilde{\theta})=S_{\psi}(\widetilde{\theta})-\mathcal{I}_{\psi\phi\cdot\gamma}(\widetilde{\theta})\mathcal{I}^{-1}_{\phi\cdot\gamma}(\widetilde{\theta})S_{\phi}(\widetilde{\theta})$ is the part of $S_{\psi}(\widetilde{\theta})$ that remains after eliminating the effect of $S_{\phi}(\widetilde{\theta})$. In summary, $S^{*}_{\psi}(\widetilde{\theta}\perp S_{\phi}(\widetilde{\theta})$, though $S_{\phi}(\widetilde{\theta})$ has “peer” effect on $S_{\psi}(\widetilde{\theta})$. Three more things regarding $RS^{*}_{\psi}(P)$ are worth noting. First, $RS^{*}_{\psi}(P)$ requires estimation only under the joint null, namely for the constrained model in which both $\psi=0$ and $\phi=0$. Given the full specification of the model $l(\gamma,\psi,\phi)$, it is of course possible to derive a RS test for $H_{0}:\psi=0$ in the presence of $\phi$. However, that requires the MLE of $\phi$, which could be difficult to obtain in some cases. Second, when $\mathcal{I}_{\psi\phi\cdot\gamma}=0$, $RS^{*}_{\psi}(P)=RS_{\psi}$. This is a simple condition to check in practice. As mentioned earlier, if this condition is true, $RS_{\psi}$ is an asymptotically valid test in the local presence of $\phi$. Finally, Bera and Yoon, (1993) showed that for local misspecification $RS^{*}_{\psi}(P)$ is asymptotically equivalent to Neyman, (1959) $C(\alpha)$ test, and therefore, shares its optimality properties.
Example 5.4 To illustrate the usefulness of the robust score statistic $RS^{*}_{\psi}(P)$, we now consider the tests developed in Anselin et al., (1996) for the mixed regressive - spatial autoregressive (SAR) model with a SAR disturbance
@@
where $\widetilde{u}=y-X\widetilde{\gamma}$ are the OLS residuals, $\widetilde{\sigma}^{2}={\widetilde{u}^{\prime}\widetilde{u}}/{n}$ and $T=tr[(W^{\prime}+W)W]$. One very interesting observation here is that $RS_{\psi}$ is essentially same as the widely used Moran, (1948) I test. Let us now consider testing $H_{0}$ under the local presence of $\phi$. First, the crucial quantity to consider is $\mathcal{I}_{\psi\phi\cdot\gamma}$ which is equal to $T$ and that can never be zero. Therefore robustification of $RS_{\psi}$ is needed. Anselin et al., (1996) derived the robust test as
@@
In a similar way we can find $RS_{\phi}$ and $RS^{*}_{\phi}(P)$ which are given, respectively, by
@@
This statistic is asymptotically distributed $\chi^{2}_{2}$. It is easy to verify that [see Bera et al., (2020, Corollary 1)]
@@
Anselin and Florax, (1995) and Anselin et al., (1996) provided simulation results on the finite sample performance of the unadjusted and adjusted RS tests and some related tests. The adjusted tests $RS^{*}_{\psi}(P)$ and $RS^{*}_{\phi}(P)$ performed remarkably well. Those had very reasonable empirical sizes, remaining within the confidence intervals in all cases. In terms of power they performed exactly the way they were supposed to.
@@
Now we combine the results of Sections 5.1 and 5.2 and develop robust tests $RS^{*}_{\psi}(DP)$ which provides a two-way protection against both types of misspecifications, distributional $(D)$ and parametric $(P)$. As we have noted in (59), $RS^{*}_{\psi}(D)$ involves both the $J(\theta)$ and $K(\theta)$ matrices in the variance expression of $S_{\psi}(\tilde{\theta})$. While to account of the parametric misspecification, as we did in (71), $\mathcal{I}_{\psi\phi\cdot\gamma}(\widetilde{\theta})\mathcal{I}^{-1}_{\phi\cdot\gamma}(\widetilde{\theta})S_{\phi}(\widetilde{\theta})$ must be subtracted from $S_{\psi}(\tilde{\theta})$ to center its mean to zero. The expression for $RS^{*}_{\psi}(DP)$ is given by [for details see Bera et al., (2020)]:
@@
We started this survey paper by stating that C.R. Rao’s work was always inspired by some practical problems. In his 2003 Econometric Theory (ET) Interview [see Bera, (2003, p.349)], on the RS test, Rao had the following to say, “The test evolved in a natural way while I was analyzing some genetic data. As I recall, the problem was the estimation of a linkage parameter using data sets from different experiments designed in such a way that each data set had information on the same linkage parameter. It was, however, necessary to test whether such an assumption could be made because of unforeseen factors affecting the experiments. This required a test for consistency of estimates derived from different experimental data sets.” Thus we had a new statistical test principle, after LR and W, motivated by a practical problem in genetics. However, as we have narrated here, the resulting RS test principle has a far reaching influence even beyond the Statistics, and in particular, it has become one of the most useful model misspecification testing tools in the Econometrics literature. In the history of any scientific field, once in a while there comes a moment for a major breakthrough. Appearance of Rao, (1948) was such a historical moment. In fact, after that we have not witnessed any new test principle, beyond the trinity, LR, W and RS.
@@ Acknowledgements
We are profoundly thankful to an anonymous referee for her/his careful reading and offering many pertinent comments which led to improvement of the paper. We are grateful to our research assistant (RA) Anirudh Adhikary. Without his assistance this paper wouldn’t have taken off. We are also grateful to our other RAs, Rong Yuwen, Tiancheng Guo, Scarlett He, Yice Zhang and Wenqi Zeng for their painstaking work on compiling the citation numbers and a very careful analysis. We wish we could report all the results they obtained. Thanks are also due to Dr. Malabika Koley for comments on an earlier draft, that improved the exposition of the paper. We however, retain the responsibility of any remaining errors. Part of this work was completed when the second author visited University of Illinois at Urbana-Champaign (UIUC).